find the fourth term of a geometric sequence in which a1=-3 and r=7
Answer by praseenakos@yahoo.com(507)

QUESTION:

find the fourth term of a geometric sequence in which a1=-3 and r=7


ANSWER:

A series of the form,
a1, a1r^1, a1r^2, a1r^3, a1r^4,........... is called a geometric progression.
where a is the first term and r is the common ratio.

nth term of a geometric progression is given by the formula,

n th term = a1r^(n-1)

here a1 = -3 and r = 7


so 4 th term = -3 * (7)^(4-1)


==> = -3 * 7 ^ 3


==> = -3*343

==> = -1029


So 4 th term = -1029


Hope you found the explanation useful.


Regards.

Praseena.